OpenStudy (oliviagggg):

What is the slope of a line that passes through the points (-2,4) and (-6,12) -2 -1/2 1/2 2

OpenStudy (stuck-help):

could you draw the line on a graph

pooja195 (pooja195):

You don't need a graph, use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \]

OpenStudy (oliviagggg):

There isn't one but u can pull it up if you type the points on math way.com

OpenStudy (oliviagggg):

There's not a pic though

pooja195 (pooja195):

Use the slope formula i gave

pooja195 (pooja195):

\[\huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope}\]

Zeronknight (zeronknight):

\(\Large\cal\color{red}{🌟🌟Welcome~🌟🌟}\color{darkgreen}{🌟🌟To~🌟🌟}\color{gold}{🌟🌟OpenStudy!🌟🌟}\)

OpenStudy (oliviagggg):

I just got it

pooja195 (pooja195):

\[\huge~\rm~\bf\color{blue}{ \frac{12-4=? }{-6--2=? }=slope}\] 2 negs make a pos \[\huge~\rm~\bf\color{blue}{ \frac{12-4=? }{-6+2=? }=slope}\] calculate the slope

OpenStudy (oliviagggg):

I got the answer thanks :)

OpenStudy (oliviagggg):

It was -2

pooja195 (pooja195):

yw ^_^ Welcome to OpenStudy :-)

OpenStudy (oliviagggg):

Thx